OpenStudy (anonymous):

integral (2)/(x^2-x)

hartnn (hartnn):

can u do partial fractions ?

hartnn (hartnn):

(2)/(x^2-x) = (2)/(x(x-1)) = A/x + B/(x-1) can u find A and B ?

OpenStudy (anonymous):

would it turn out to be -2ln(x)+2ln(x-1)+c?

hartnn (hartnn):

yes, thats correct.

OpenStudy (anonymous):

woo thank you

hartnn (hartnn):

\[ \begin{array}l\color{red}{\text{w}}\color{orange}{\text{e}}\color{#e6e600}{\text{l}}\color{green}{\text{c}}\color{blue}{\text{o}}\color{purple}{\text{m}}\color{purple}{\text{e}}\color{red}{\text{ }}\color{orange}{\text{^}}\color{#e6e600}{\text{_}}\color{green}{\text{^}}\color{blue}{\text{}}\end{array} \]